OpenStudy (anonymous):

Susie wants to make a quilt in the shape of a golden rectangle. She only has enough binding material for a quilt with a perimeter of 31 feet. What should the dimensions of the quilt be?

OpenStudy (anonymous):

A golden rectangle has the following dimensions:  which satisfy \[\frac{a+b}{a}=\frac{a}{b}\approx1.618\] The perimeter of such a rectangle will be \[P=2a+2(a+b)=4a+2b\] Cross-multiplying in the ratio equation, you get \[(a+b)b=a^2~~\iff~~a^2-ab-b^2=0\] Solve for one of the variables in the perimeter equation: \[4a+2b=31~~\iff~~\begin{cases}a=\dfrac{31-2b}{4}&\text{or}\\\\ b=\dfrac{31-4a}{2}\end{cases}\] then substitute into the modified ratio equation: \[\left(\frac{31-2b}{4}\right)^2-\left(\frac{31-2b}{4}\right)b-b^2=0\\~~~~~~~~~~~~~~~~~~~~~~~~\text{or}~~\\ a^2-a\left(\frac{31-4a}{2}\right)-\left(\frac{31-4a}{2}\right)^2=0\] Whichever path you take the answer will be the same.